OpenStudy (anonymous):

So i'm practicing for my ACT compass test and i've come across this questions Saying that 4 < √x < 9 is equivalent to saying what about x ? A. 0 < x < 5 B. 0 < x < 65 C. 2 < x < 3 D. 4 < x < 9 E. 16 < x < 81 Now after researching everyone is giving me answer E and i understand that... BUT.. The answer sheet says its B... WHY! I don't understand @.@

OpenStudy (mandre):

Take the square of each term 4 < √x < 9 4^2 = 16 9^2=81 (√x ) = ?

OpenStudy (anonymous):

Maybe its a misprint?

OpenStudy (mandre):

Last one should be (√x )^2 = ?

OpenStudy (mandre):

Answer sheet is wrong.

OpenStudy (anonymous):

Bingo.

OpenStudy (anonymous):

You think?.. UG.. stupid test -_-

OpenStudy (anonymous):

Draw the graph and see....Clearly, when sqrt(x) is between 3 and 4, then we get 9 < x < 16. You could also just square each part of the inequality to get the same solution. Note that we discard our negative x-values since the co-domain of sqrt(x) is restricted and hence we only look at the principal root and ignore the negative roots.